# UniHash: Unifying Pointwise and Pairwise Hashing Paradigms for Seen and Unseen Category Retrieval

### v2

@@ 4.2 Generalization to Unseen Categories
To evaluate retrieval generalization on unseen classes, following [30], we split datasets into seen and unseen categories, holding out 20% of CIFAR-10 and 15% of ImageNet and MSCOCO as unseen and excluding them from training. Table 2 reports mAP for four cases: (a) Seen@Seen and Seen@All: queries from seen classes searching among databases of seen-only or all images, and (b) Unseen@Unseen and Unseen@All: queries from novel classes searching among only unseen images or the entire database. It is worth noting that this constitutes a more challenging zero-shot hashing task [30]. Although zero-shot hashing has been studied in prior works [37, 5], their experimental settings typically designate only a single category as unseen during testing, which provides a limited assessment of a method’s retrieval capability when facing multiple unseen categories.
As shown in Table 2, UniHash demonstrates robust performance on both seen and unseen retrieval tasks. For seen-category queries, UniHash achieves near-perfect accuracy (e.g.,  97% mAP on CIFAR-10) and sustains this on the Seen@All task. Competing methods also perform well but often drop slightly on Seen@All due to interference from unseen classes. More importantly, UniHash significantly outperforms prior methods on unseen-category queries. On CIFAR-10, UniHash achieves 84.7% mAP on Unseen@Unseen (vs. 82.0% for OrthoHash) and maintains a much higher mAP on Unseen@All (35.3% vs. 7.2% for CSQ). Similar trends hold on ImageNet and MSCOCO, where UniHash yields the highest Unseen@Unseen accuracy and clearly better Unseen@All mAP (with about 4% absolute gains over the best baselines). These results indicate that UniHash enables strong generalization, retrieving novel-class images even with many distractors.
@@ 4.3 Ablation Study
Analysis of Overall Framework. To evaluate the contribution of each component in overall architecture, we conducted a closed-set ablation study on both the SM-MoH and Mutual Learning (ML) loss modules, as shown in Table 3. Baseline (without either SM-MoH or ML) shows that the center-based branch consistently outperforms the pairwise branch (e.g., 0.8940 vs. 0.8873 on ImageNet at 64 bits), indicating stronger standalone effectiveness. ML alone improves the weaker pairwise branch (e.g., 0.8894 vs. 0.8873 pairwise branch on ImageNet at 64 bits), by enabling bidirectional knowledge transfer. Gains for the center-based branch are smaller due to limited diversity without SM-MoH. SM-MoH alone slightly reduces performance (e.g., center-based: 0.8325 vs. 0.8940), suggesting that without collaborative training, the structured interaction it introduces is underutilized. SM-MoH + ML achieves the best results (e.g., 0.8997/0.9062 on ImageNet at 64 bits), as SM-MoH enhances inter-branch communication and mutual learning amplifies mutual supervision, improving both intra-class compactness and inter-class separability.
@@
Hyperparameter Tuning. Figure 6 shows a hyperparameter study on ImageNet with 16-bit hash codes, varying one of $\lambda_{1}$, $\lambda_{2}$, or $\lambda_{3}$ while fixing the others. $\lambda_{1}$ and $\lambda_{2}$ control the center- and pairwise-based losses, and $\lambda_{3}$ regulates mutual loss. The results reveal a dominant–auxiliary dynamic: when $\lambda_{1}\gg\lambda_{2}$, the center branch outperforms; conversely, when $\lambda_{2}\gg\lambda_{1}$, the pairwise branch outperforms. Configurations with $\lambda_{1}>\lambda_{2}$ achieve higher mAP, so the center branch serves as the primary while the pairwise branch fine-tunes it. Both $\lambda_{1}$ and $\lambda_{2}$ exhibit unimodal trends, with optimal values within $[1,10]$. And $\lambda_{3}$ performs best at 11, beyond which excessive coupling degrades performance. The optimal setting is $\lambda_{1}=4$, $\lambda_{2}=1$, $\lambda_{3}=1$.
(a) $\lambda_{1}$ Tuning Effects
(b) $\lambda_{2}$ Tuning Effects
(c) $\lambda_{3}$ Tuning Effects
@@ 4.4 Hash Codes Visualization
To illustrate the effectiveness of UniHash, we visualize the t-SNE [32] of hash codes generated by three methods on ImageNet, as shown in Fig. 5: (a) center-based [34], (b) pairwise-based [47], and (c) our UniHash approach. The center-based method in (a) produces a circular distribution, indicating clearly decent global alignment but limited fine-grained separation. The pairwise-based method in (b) yields a more elongated structure with overlapping clusters, suggesting weaker overall structure. In contrast, UniHash in (c) consistently leads to a more well-clustered distribution, with fewer ambiguous points across class boundaries. These results show that UniHash improves intra-class compactness and inter-class separability by making the center- and pairwise-based branches complementary, yielding more discriminative hash codes.
@@ Appendix A Summary
Appendix B Training Setup
B.1 Datasets
ImageNet is a large-scale image classification dataset consisting of over 1.2 million images annotated with 1,000 categories. Following the protocol in [9], we use the ILSVRC2012 version for evaluation. The validation set of 50K images is used as the query set, while the remaining training images form the database. For training, we randomly sample 130K images from the database.
CIFAR-10 consists of 60,000 images across 10 categories, with each image sized $32\times 32$. Following the standard practice in [19], we use 10K test images as the query set and the remaining 50K training images as the database. For training, 5K images are randomly sampled from the database.
MSCOCO  [24] is an image recognition, segmentation, and captioning dataset. We use the public version processed by [24], where images with missing category information have been filtered out. This results in 122K labeled images by combining the training and validation splits. We randomly sample 5K images as the query set, with the remaining images forming the database, and then randomly sample 10K images from the database for training.
Seen/Unseen Splitting [30] refers to our unified protocol for constructing seen and unseen subsets across both single-label and multi-label datasets. For single-label datasets such as CIFAR-10 and ImageNet-1K, we partition the label space using dataset-specific ratios: in CIFAR-10, 80% of the categories are designated as seen, whereas in ImageNet, 85% of the categories are treated as seen; the remaining categories in each dataset serve as unseen. For the multi-label dataset MSCOCO, we likewise divide the label space by a category ratio (85% seen). However, since images may contain multiple tags, an image is included in the seen or unseen subset only when all of its labels fall entirely within the corresponding class group, while images with mixed labels are used exclusively in the “all-class” database settings. Using these purified subsets, we then construct the four evaluation protocols—Seen@Seen, Seen@All, Unseen@Unseen, and Unseen@All—in a consistent manner across all datasets.
@@ B.1 Datasets
ImageNet is released under a non-commercial license, and the use of the dataset is restricted to research and educational purposes. Users must apply for access and agree to the ImageNet Terms of Use. CIFAR-10 is made publicly available by the University of Toronto under the MIT License. This permits free use, modification, and distribution of the dataset for both research and commercial purposes. For MSCOCO, the annotations are provided under the Creative Commons Attribution 4.0 License ($C$ BY 4.0), and the use of the images must comply with the Flickr Terms of Use. The dataset is released for academic and research use.
Appendix C Ablation Study and Further Analysis
C.1 Comparison with Traditional Mutual Learning
Traditional Deep Mutual Learning (DML) [44, 39, 46, 45] typically employs two separate branches, where each branch contains an independently initialized and trained deep neural network (DNN). While this design allows for mutual supervision between diverse learners, it also limits the potential for fine-grained interaction between the learned representations, especially in the context of hashing where compact and consistent binary codes are desired.
@@ C.2 SM-MoH Module Analysis
Design of Hashing Experts vs. Traditional Experts. Traditional Mixture of Experts (MoE) [6, 7, 28, 29] typically employs two-layer MLPs with ReLU activations as experts, designed for general-purpose representation transformation. In contrast, our MoH replaces these with specialized hashing experts, which directly map the feature dimension to the hash bit dimension — effectively acting as task-specific hashing layers.
@@ C.3 SM-MoH Parameter Tuning
Appendix D Theory Supplement
@@
D.1 Risk Comparison: Baseline vs. UniHash
@@
D.2 Seen/Unseen Generalization
@@
Let $f_{\text{base}}$ be a single-branch optimum and $f_{\text{UniHash}}$ the UniHash solution. Combining Sec. D.1 with the two effects above yields
@@
Appendix E Limitations
While our proposed Unified Hashing (UniHash) consistently demonstrates strong performance across diverse datasets and settings, several complementary aspects remain open for further exploration. First, the dual-branch architecture and the SM-MoH module introduce a more intricate interaction mechanism in the hash code generation process, which makes the interpretability of individual bits less straightforward. Second, like many deep hashing methods, UniHash may encounter practical limitations in resource-constrained environments due to its reliance on deep neural networks. However, as numerous methods exist to improve the deployability of deep neural networks, this limitation has been less significant. These considerations highlight valuable directions for future work while underscoring the effectiveness of our method.

### v3

@@ 4.2 Generalization to Unseen Categories
To evaluate retrieval generalization on unseen classes, following (Shen et al., 2025), we split datasets into seen and unseen categories, holding out 20% of CIFAR-10 and 15% of ImageNet and MSCOCO as unseen and excluding them from training. Table 2 reports mAP for four cases: (a) Seen@Seen and Seen@All: queries from seen classes searching among databases of seen-only or all images, and (b) Unseen@Unseen and Unseen@All: queries from novel classes searching among only unseen images or the entire database. It is worth noting that this constitutes a more challenging zero-shot hashing task (Shen et al., 2025). Although zero-shot hashing has been studied in prior works (Wei et al., 2023; Cao et al., 2017b), their experimental settings typically designate only a single category as unseen during testing, which provides a limited assessment of a method’s retrieval capability when facing multiple unseen categories.
As shown in Table 2, UniHash demonstrates robust performance on both seen and unseen retrieval tasks. For seen-category queries, UniHash achieves near-perfect accuracy (e.g.,  97% mAP on CIFAR-10) and sustains this on the Seen@All task. Competing methods also perform well but often drop slightly on Seen@All due to interference from unseen classes. More importantly, UniHash significantly outperforms prior methods on unseen-category queries. On CIFAR-10, UniHash achieves 84.7% mAP on Unseen@Unseen (vs. 82.0% for OrthoHash) and maintains a much higher mAP on Unseen@All (82.3% vs. 7.2% for CSQ). Similar trends hold on ImageNet and MSCOCO, where UniHash yields the highest Unseen@Unseen accuracy and clearly better Unseen@All mAP (with about 4% absolute gains over the best baselines). These results indicate that UniHash enables strong generalization, retrieving novel-class images even with many distractors.
@@ 4.3 Ablation Study
Analysis of Overall Framework. To evaluate the contribution of each component in overall architecture, we conducted a closed-set ablation study on both the SM-MoH and Mutual Learning (ML) loss modules, as shown in Table 3. Baseline (without either SM-MoH or ML) shows that the center-based branch consistently outperforms the pairwise branch (e.g., 0.8940 vs. 0.8873 on ImageNet at 64 bits), indicating stronger standalone effectiveness. ML alone improves the weaker pairwise branch (e.g., 0.8894 vs. 0.8873 pairwise branch on ImageNet at 64 bits), by enabling bidirectional knowledge transfer. Gains for branches are smaller due to limited diversity without SM-MoH. SM-MoH is not a stand-alone architectural add-on; it is designed to facilitate cross-branch structured interaction. In isolation, SM-MoH slightly reduces performance (e.g., center-based: 0.8325 vs. 0.8940), whereas combining SM-MoH with mutual learning yields the best results (e.g., 0.8997/0.9062 on ImageNet at 64 bits). This trend suggests that SM-MoH strengthens inter-branch communication, and mutual learning further amplifies complementary supervision, jointly improving intra-class compactness and inter-class separability.
@@
(a) $\lambda_{1}$ Tuning Effects
(b) $\lambda_{2}$ Tuning Effects
(c) $\lambda_{3}$ Tuning Effects
@@
Hyperparameter Tuning. Figure 5 shows a hyperparameter study on ImageNet with 16-bit hash codes, varying one of $\lambda_{1}$, $\lambda_{2}$, or $\lambda_{3}$ while fixing the others. $\lambda_{1}$ and $\lambda_{2}$ control the center- and pairwise-based losses, and $\lambda_{3}$ regulates mutual loss. The results reveal a dominant–auxiliary dynamic: when $\lambda_{1}\gg\lambda_{2}$, the center branch outperforms; conversely, when $\lambda_{2}\gg\lambda_{1}$, the pairwise branch outperforms. Configurations with $\lambda_{1}>\lambda_{2}$ achieve higher mAP, so the center branch serves as the primary while the pairwise branch fine-tunes it. Both $\lambda_{1}$ and $\lambda_{2}$ exhibit unimodal trends, with optimal values within $[1,10]$. And $\lambda_{3}$ performs best at $1$, beyond which excessive coupling degrades performance. The optimal setting is $\lambda_{1}=4$, $\lambda_{2}=1$, $\lambda_{3}=1$.
@@ 4.4 Hash Codes Visualization
To illustrate UniHash, we visualize the t-SNE (van der Maaten and Hinton, 2008) of hash codes generated by three methods on ImageNet, as shown in Fig. 4: (a) center-based (Wang et al., 2023), (b) pairwise-based (Zheng et al., 2020), and (c) our UniHash approach. The center-based method in (a) produces a circular distribution, indicating decent global alignment but limited fine-grained separation. The pairwise-based method in (b) yields a more elongated structure with overlapping clusters, suggesting weaker overall structure. In contrast, UniHash in (c) consistently leads to a more well-clustered distribution, with fewer ambiguous points across class boundaries. These results show that UniHash improves intra-class compactness and inter-class separability by making the center- and pairwise-based branches complementary, yielding more discriminative hash codes.
@@ Appendix A Summary
Appendix B Closed-Set Training Performance
(a) 16 bit PR Curves
(b) 32 bit PR Curves
(c) 64 bit PR Curves
Appendix C Training Setup
C.1 Datasets
ImageNet is a large-scale image classification dataset consisting of over 1.2 million images annotated with 1,000 categories. Following the protocol in [Deng et al., 2009], we use the ILSVRC2012 version for evaluation. The validation set of 50K images is used as the query set, while the remaining training images form the database. For training, we randomly sample 130K images from the database.
CIFAR-10 consists of 60,000 images across 10 categories, with each image sized $32\times 32$. Following the standard practice in [Krizhevsky and Hinton, 2009], we use 10K test images as the query set and the remaining 50K training images as the database. For training, 5K images are randomly sampled from the database.
MSCOCO  [Lin et al., 2014] is an image recognition, segmentation, and captioning dataset. We use the public version processed by [Lin et al., 2014], where images with missing category information have been filtered out. This results in 122K labeled images by combining the training and validation splits. We randomly sample 5K images as the query set, with the remaining images forming the database, and then randomly sample 10K images from the database for training.
Seen/Unseen Splitting [Shen et al., 2025] refers to our unified protocol for constructing seen and unseen subsets across both single-label and multi-label datasets. For single-label datasets such as CIFAR-10 and ImageNet-1K, we partition the label space using dataset-specific ratios: in CIFAR-10, 80% of the categories are designated as seen, whereas in ImageNet, 85% of the categories are treated as seen; the remaining categories in each dataset serve as unseen. For the multi-label dataset MSCOCO, we likewise divide the label space by a category ratio (85% seen). However, since images may contain multiple tags, an image is included in the seen or unseen subset only when all of its labels fall entirely within the corresponding class group, while images with mixed labels are used exclusively in the “all-class” database settings. Using these purified subsets, we then construct the four evaluation protocols—Seen@Seen, Seen@All, Unseen@Unseen, and Unseen@All—in a consistent manner across all datasets.
@@ C.1 Datasets
ImageNet is released under a non-commercial license, and the use of the dataset is restricted to research and educational purposes. Users must apply for access and agree to the ImageNet Terms of Use. CIFAR-10 is made publicly available by the University of Toronto under the MIT License. This permits free use, modification, and distribution of the dataset for both research and commercial purposes. For MSCOCO, the annotations are provided under the Creative Commons Attribution 4.0 License (CC BY 4.0), and the use of the images must comply with the Flickr Terms of Use. The dataset is released for academic and research use.
Appendix D Ablation Study and Further Analysis
D.1 Comparison with Traditional Mutual Learning
Traditional Deep Mutual Learning (DML) [Zhang et al., 2018, Wu et al., 2019, Zhao et al., 2021, Zhao and Han, 2021] typically employs two separate branches, where each branch contains an independently initialized and trained deep neural network (DNN). While this design allows for mutual supervision between diverse learners, it also limits the potential for fine-grained interaction between the learned representations, especially in the context of hashing where compact and consistent binary codes are desired.
@@ D.2 SM-MoH Module Analysis
Design of Hashing Experts vs. Traditional Experts. Traditional Mixture of Experts (MoE) [Chen et al., 2023, 2022, Riquelme et al., 2021, Shazeer et al., 2017] typically employs two-layer MLPs with ReLU activations as experts, designed for general-purpose representation transformation. In contrast, our MoH replaces these with specialized hashing experts, which directly map the feature dimension to the hash bit dimension — effectively acting as task-specific hashing layers.
@@ D.3 SM-MoH Parameter Tuning
D.4 Effectiveness of SM-MoH Design
SM-MoH is not introduced to boost any individual branch in isolation. Rather, it is specifically designed to facilitate mutual learning by providing a shared expert space and adaptive routing that enhance cross-branch knowledge transfer and alignment between the pointwise and pairwise streams.
Why using SM-MoH alone can lead to a performance drop.
The reason is that center-based supervision essentially pulls samples toward their class/hash centers, yielding a signal that is relatively global, stable, and low-variance. In contrast, SM-MoH introduces (i) dynamic routing of gating with top-$k$ experts, meaning that samples from the same class may be dispatched to different experts, and (ii) expert specialization, where different experts learn different mappings. When only the center loss is present, there is insufficient constraint to coordinate the outputs across experts. Accordingly, when SM-MoH is used without mutual learning, it may not provide benefits and can even underperform the plain center-based baseline (Table 3).
Why mutual learning alone is limited.
Mutual learning without SM-MoH brings only a small gain (Table 3), because pointwise and pairwise supervision induce representations with different geometric biases, so the two branches often lie in partially misaligned subspaces, limiting effective cross-branch knowledge transfer.
Why the combination works.
When SM-MoH combined with mutual learning, the pairwise branch provides relational constraints among samples, offering a finer-grained and stronger structural signal, while the center-based branch contributes complementary global guidance as shown in Table 3. Mutual learning further pulls the two branches into a shared hash space, effectively acting as a cross-branch consistency regularizer for the expert routing and outputs. As a result, experts are jointly constrained by the pairwise relational structure and cross-branch alignment, enabling SM-MoH to capture diverse modes/hard cases via multiple experts without sacrificing intra-class consistency.
Therefore, we emphasize that SM-MoH is an interaction-centric component for cross-branch integration, rather than a single-branch booster.
Appendix E Theory Supplement
@@
E.1 Risk Comparison: Baseline vs. UniHash
@@
E.2 Seen/Unseen Generalization
@@
Let $f_{\text{base}}$ be a single-branch optimum and $f_{\text{UniHash}}$ the UniHash solution. Combining Sec. E.1 with the two effects above yields
@@
Appendix F Computational and Space Complexity
We analyze UniHash from both computational and space complexity perspectives. As shown below, inference scales linearly with the batch size and relies on compact binary codes, so the retrieval pipeline remains efficient in practice; in turn, this efficiency offsets the extra complexity introduced by our novel structure, making UniHash suitable for large-scale retrieval.
F.1 Computational Complexity
The overall computational complexity is $\mathcal{O}\!\left(N\left(d^{2}+kdq\right)\right)$, where $N$ denotes the batch size, $d$ denotes the output dimensionality of the backbone, $q$ is the length of the binary hash code, and $k$ is the number of experts selected per sample.
A detailed analysis is presented below. Each gating network consists of two MLP layers, resulting in a computational complexity of $\mathcal{O}(d^{2})$, where $d$ denotes the output dimensionality of the backbone. Similarly, each hash expert is composed of two MLP layers, leading to a complexity of $\mathcal{O}(dq)$, where $q$ is the length of the binary hash code. For each sample, the complexity is $\mathcal{O}(d^{2}+kdq)$, where $k$ is the number of experts selected per sample. Therefore, the overall computational complexity is $\mathcal{O}\!\left(N\left(d^{2}+kdq\right)\right)$, where $N$ denotes the batch size.
F.2 Space Complexity
The overall space complexity is $\mathcal{O}\!\left(d+m(d+q)\right)$, where $d$ denotes the output dimensionality of the backbone, $q$ is the length of the binary hash code, and $m$ is the total number of experts.
A detailed analysis is presented below. Each gating network consists of two MLP layers, resulting in a space complexity of $\mathcal{O}(d)$, where $d$ denotes the output dimensionality of the backbone. Similarly, each hash expert is composed of two MLP layers, leading to a complexity of $\mathcal{O}(d+q)$, where $q$ is the length of the binary hash code. Since there are $m$ experts, the overall space complexity is $\mathcal{O}\!\left(d+m(d+q)\right)$.
Appendix G Limitations
Is this work merely an incremental combination of existing ingredients, rather than a genuinely new contribution?
Our key clarification is that the paper is not optimizing a single paradigm with add-ons, but explicitly targets the structural tension between seen-category accuracy and unseen-category generalization, and proposes a unified dual-branch training formulation to resolve it. In this view, the dual-branch design and the alignment objective are not optional embellishments: they operationalize the complementary failure modes of pointwise vs. pairwise supervision under the same retrieval goal, which is different from prior works that stay within one paradigm and only refine losses or architectures.
Does SM-MoH really contribute on its own, does it only work when paired with mutual learning?
A strong concern is that SM-MoH alone can degrade performance (e.g., the center-based branch drops notably when SM-MoH is inserted without mutual learning), which could be interpreted as lack of standalone utility or robustness. The paper’s intended claim, however, is that SM-MoH is not designed as a universally beneficial plug-in; it is a structural mechanism to enable cross-paradigm interaction—routing and expert specialization create diversity, while the mutual-learning alignment provides the missing constraint that prevents intra-class dispersion across experts. Therefore, the relevant evaluation is the coupled system: SM-MoH + mutual learning, where ablations show the drop disappears and performance peaks, supporting that the module’s benefit is conditional but purposeful.
Is the practical efficiency worth the added complexity?
Our design is simplifiable at inference: retrieval only relies on the binary-code generation path, and we can follow the paper’s select-the-better-branch strategy to keep only a single branch for encoding both queries and the database, making inference cost comparable to standard single-branch methods. Moreover, SM-MoH activates only a sparse set of experts via top-k routing, which keeps its marginal compute controlled. The complexity analysis also indicates that most additional overhead comes from the dual-branch interaction during training, which is an offline one-time cost and does not affect online retrieval throughput. Therefore, UniHash remains suitable for large-scale retrieval.